OpenStudy (anonymous):

The equation of the line through (8,6) and (2,-4) is a.)5x-3y=22 b.)y=3/5x+8/5 c.) 3x+4y=48 d.) -4x+2y=-16

OpenStudy (anonymous):

y - y1 = m(x - x1) y-6=5/3(x-8) y-6=5/3x-10 y=5/3x-4 how would I simplify it?

Nnesha (nnesha):

b is not -4

Nnesha (nnesha):

slope is right 5/3 try to find b again :)

OpenStudy (anonymous):

It is -40/3 plus 6=-22/3 ?

OpenStudy (anonymous):

would it be a?

Nnesha (nnesha):

yes it is a

Nnesha (nnesha):

but when we solve first equation for y we cannot get -4

Nnesha (nnesha):

 this is how you should solve for b